The function F described by F(C) =
95
C + 32 gives the Fahrenheit temperature corresponding to the
Celsius temperature C. Find the Fahrenheit temperature equivalent to -25 °C.

Answer by stanbon(75887)

The function F described by F(C) = (9/5)C + 32
gives the Fahrenheit temperature corresponding to the
Celsius temperature C.
Find the Fahrenheit temperature equivalent to -25 °C.
----
F(-25) = (9/5)(-25) + 32
---
F(-25) = -45+32 = -13 degrees Fahrenheit
==============================================
Cheers,
Stan H.
===========